# Matrix-free Second-order Optimization of Gaussian Splats with Residual Sampling

### v2

@@ 4.4 Estimation of Loss Function with Residual Sampling
$L=\int_{\Omega}r(\mathbf{p})^{2}\,d\mathbf{p}=\int_{\Omega}\ell(\mathbf{p})\,d\mathbf{p},$ (11)
where the integral is defined over $\Omega$, the union of all image domains of all cameras, and calculated at the coordinates $\mathbf{p}\in\mathbb{R}^{2}$ of the image plane. We also define $\ell(\mathbf{p})=r(\mathbf{p})^{2}$ to remove the squared term, which simplifies the subsequent derivations. To estimate the loss function $L$, we express it as an expectation with respect to an arbitrary probability distribution $q(\mathbf{p})$:
$L=\int_{\Omega}\ell(\mathbf{p})\frac{q(\mathbf{p})}{q(\mathbf{p})}\,d\mathbf{p}=\mathbb{E}_{\mathbf{p}\sim q(\mathbf{p})}\left[\frac{\ell(\mathbf{p})}{q(\mathbf{p})}\right].$ (12)
@@
$L\approx\hat{L}=\frac{1}{N}\sum_{i=1}^{N}\ell(i)\,\frac{1}{q(i)}.$ (13)
@@
$\frac{\partial\hat{L}}{\partial\boldsymbol{\beta}_{j}}=\frac{1}{N}\sum_{i=1}^{N}\frac{\partial\ell(i)}{\partial\boldsymbol{\beta}_{j}}\,\frac{1}{q(i)}=\frac{2}{N}\sum_{i=1}^{N}\frac{\partial r_{i}}{\partial\beta_{j}}r_{i}\,\frac{1}{q(i)}$ (14)
This gradient corresponds to the right-hand side of the normal equation given in Eq. 8 since $\frac{\partial r_{i}}{\partial\beta_{j}}=J_{ij}$.
@@
4.5 Dynamic Learning Rate Scheduler
When second-order optimizers are applied to large-scale systems, numerical errors can occur, causing the solution vector $\Delta\boldsymbol{\beta}$ to overshoot the true loss landscape. To determine the optimal learning rate, line search algorithms or trust region methods [armijo1966minimization, hollein20243dgs, more1994line, conn2000trust, nocedal1999numerical, bottou2018optimization] are usually employed. However, these methods require additional forward or backward passes, increasing the computational overhead and slowing down the overall algorithm. Rather than relying on these methods, we estimate the learning rate by constraining the maximum update of color values. In 3DGS, the color parameters range between $\approx-1.77$ and $\approx 1.77$ due to the level 0 spherical harmonic coefficient. This gives us a natural bound for the color parameters. We trust the update direction $\Delta\boldsymbol{\beta}$, as long as the resulting color change does not exceed 1. If the change surpasses this threshold, we scale $\Delta\boldsymbol{\beta}$ so that the maximum change remains 1. Notably, this scaling is applied uniformly across all parameters, regardless of their type (e.g., opacity, color, mean, scale, or rotation). The effectiveness of this learning rate heuristic is demonstrated in Tab. 4, and examples of assigned learning rates are given in the Supplementary.
5 Results
Datasets and Metrics. We conduct our main experiments on the synthetic NeRF [nerf] and real-world Mip-NeRF360 indoor datasets [barron2021mipnerf]. In the synthetic dataset, we position 10,00010{,}000 Gaussians at random locations with random colors inside of the cube encapsulating the object. In the real-world scenes, the Gaussians are initialized with structure-from-motion point clouds as in 3DGS [3dgs]. For all experiments, we use the default training and test splits. We report the test set performance based on structural similarity index (SSIM) [ssim], learned perceptual image patch similarity score (LPIPS) [lpips], and peak signal-to-noise ratio (PSNR).
Baselines. In the synthetic dataset, we compare our optimizer against Adam [adam], RMSprop [rmsprop], and SGD with momentum [sgd_momentum], using their respective implementations in PyTorch. The gradients are computed using either the vanilla 3DGS or Taming 3DGS [taming], which enhances the efficiency of the backward computation by storing the gradient state at every $32^{\text{nd}}$ splat.
In the real-world scenes, we compare against vanilla 3DGS [3dgs] and 3DGS-LM [hollein20243dgs]. Since 3DGS-LM’s code is not publicly available, we contacted the authors, who kindly provided PSNR scores, runtime, and average memory usage across several datasets. In addition, they provided results obtained without the Adam optimizer, which are denoted in the tables as 3DGS-LM (w/o Adam). First-order baselines are run for $10,000$ iterations, and we performed a hyperparameter search to find the best learning rates.
Implementation Details. We conducted synthetic dataset experiments on a NVIDIA Tesla A40 GPU and real-world experiments on a single NVIDIA Tesla A100 GPU. Our method and the baselines optimize the mean squared error (MSE) loss. We set the sampled pixels per tile ($N$) to 32 in each experiment. We use a fixed damping parameter $\lambda$ of $0.1$ for the synthetic NeRF scenes and Mip-NeRF360 dataset. For the synthetic dataset, the batch size is fixed at $8$; for real-world datasets, it is initialized at $16$ and increased to $32$ after the 50th iteration. The maximum number of conjugate gradient iterations starts at $3$ for synthetic datasets and at $5$ for real-world datasets, rising to $8$ after the 50th iteration.
[…]
3DGS 0.854 28.21 252 5.81
3DGS-LM (w/o Adam) - 26.72 1186 $\approx$ 50
3DGS-LM - 28.31 398 $\approx$ 50
LM-RS (Ours) 0.859 28.59 190 14.75
@@ 5.1 Comparison
Please refer to the Supplementary for the results using SSIM loss, as well as quantitative results in Tanks & Temples [knapitsch2017tanks] and DeepBlending scenes [hedman2018deep].
@@ 5.2 Ablation Studies
Dynamic Learning Rate Scheduler. We compare our learning rate heuristic introduced in Sec. 4.5 against uniform learning rate, Armijo line search [armijo1966minimization, nocedal1999numerical], and a grid search method in Tab. 4. The grid search method selects the learning rate at each iteration by identifying the value that results in the greatest reduction in the objective function. Although this method is stable, the learning rates obtained are too pessimistic, and it cannot reach the quality of our heuristic. Armijo line search improves the grid search method by picking the highest learning rate that results in sufficient reduction of the loss. While this approach has been successfully integrated as an improvement in deterministic second-order optimizers, our findings indicate that its effectiveness is limited in stochastic settings, such as ours, where optimization relies on multiple approximations. We also found that the uniform learning rate can diverge (LR=0.1) or behave suboptimally (LR=0.07) compared to our scheduler.
@@
6 Discussion and Conclusion
Limitations. We use mean squared error instead of mean absolute error and only a diagonal approximation of the original SSIM loss, unlike the original Adam-based optimization for 3D Gaussian Splatting. This discrepancy in loss formulations may affect convergence behavior, but future work could explore re-weighting strategies of residuals to obtain different norms.
By leveraging the inherent sparsity of the Jacobian matrix and integrating a GPU-parallelized conjugate gradient solver, our method significantly reduces the computational overhead. Our novel view and pixel-wise sampling further enhance efficiency, enabling rapid convergence by decreasing the per-step overhead. Additionally, our dynamic learning rate scheduler eliminates the need for costly line search procedures, further accelerating training. Our approach achieves up to $5\times$ speedup over Adam, particularly excelling in scenarios with low number of Gaussians. Overall, our results highlight the potential of second-order methods in accelerating optimization for 3D Gaussian splatting. We anticipate that future work will refine these techniques, particularly incorporating additional loss terms, further advancing the efficiency and quality of Gaussian-based scene representations.
@@ Appendix A Details of CUDA Implementation
Algorithm 3 Automatic Forward Differentiation - LM-RS
@@
We follow the parallelization pattern applied in 3DGS [3dgs], dividing images into $16\times 16$ blocks, and launching a thread for each pixel. Our main difference from 3DGS is that our block size is significantly less than $16\times 16=256$, due to pixel sampling. In our implementation, we only support sample sizes that are multiples of $32$, which allows us to do fast warp-level reductions when needed. Now, we describe our CUDA kernels in detail:
Jv Kernel. This kernel is similar to the backward pass of 3DGS and is executed in two kernel calls using backward mode automatic differentiation. The first kernel computes and forwards the derivatives associated with splatting operations to the Gaussians, while the second kernel computes the derivatives related to Gaussian preprocessing. The main difference with 3DGS is that the number of threads per block is significantly reduced due to sampling, enabling us to perform faster computations. We use warp-level reductions to collect the partial dot products when needed.
Jv Kernel. This kernel implements forward‐mode automatic differentiation via the dual‐number data structure. We overload each operator so that, in a single pass, it performs both the real‐value computation and its derivative. This approach is more efficient than using backward-mode automatic differentiation when the Jacobian-vector product is needed [baydin2018automatic].
Conceptually, this kernel works similarly to the forward pass of 3DGS, but enhanced to carry the derivative information. Therefore, unlike Jv kernel, it first carries out Gaussian preprocessing, then implements splatting. Since we should additionally carry the derivative information, the shared memory usage is doubled compared to the forward pass of 3DGS. For a comparison between the forward pass of 3DGS and our forward-mode automatic differentiation, see Alg. 2 and Alg. 3. As it can be observed from the algorithm, there is no need for warp reductions in this kernel.
@@ Appendix B More Results in Synthetic Dataset
We compare our optimizer against Adam [adam], RMSprop [rmsprop], and SGD with momentum [sgd_momentum], using their respective implementations in PyTorch. The momentum term of SGD and second-moment factor of RMSprop is set to $0.99$. For SGD, we used the learning rates $0.16$ for the location, $0.2$ for color and $0.1$ for other parameters. We also apply a learning rate schedule for the mean value parameters, as suggested in 3DGS. The gradients are computed using either the vanilla 3DGS or Taming 3DGS [taming]. All baselines are run for 10,00010{,}000 iterations. We provide the per-scene scores obtained in the NeRF synthetic dataset [nerf] in Table 6. Please also see the supplementary video, comparing Adam and our optimizer in real-time. Additionally, we provide visual results of competing methods in Fig. 6.
Results with SSIM Loss: We also conduct experiments using SSIM loss, which is approximated diagonally, as mentioned in the limitations section of the main paper. The usual SSIM loss propagates gradients through the local neighbors, which is hard to implement with the parallelization scheme of our $\mathbf{Diag(J^{\top}J)}$ kernel. In Table 7, we show that our optimizer still outperforms the competing methods even with the diagonal approximation. However, to compensate for this discrepancy, we ran our method for $300$ iterations instead of $200$ iterations. The weights are set to $1.0$ and $0.2$, for MSE and SSIM terms respectively. All other hyperparameters remain identical to those specified in the main paper.
@@ Appendix C More Results in Real-world Scenes
Results with SSIM Loss: Both 3DGS-LM and our method employ the diagonal approximation of the SSIM loss. In Tab. 11, we compare both methods and show that our optimizer converges faster than 3DGS-LM (w/o Adam). However, in this setting, the diagonal approximation becomes overly restrictive, and 3DGS-LM outperforms our method by leveraging the full SSIM loss thanks to the Adam optimizer.

### v3

@@ 4.4 Estimation of Loss Function with Residual Sampling
$L=\int_{\Omega}r(\mathbf{p})^{2}\,d\mathbf{p}$ (11)
where the integral is defined over $\Omega$, the union of all image domains of all cameras, and calculated at the coordinates $\mathbf{p}\in\mathbb{R}^{2}$ of the image plane. To estimate the loss function $L$, we express it as an expectation with respect to an arbitrary probability distribution $q(\mathbf{p})$:
$L=\int_{\Omega}r(\mathbf{p})^{2}\,\frac{q(\mathbf{p})}{q(\mathbf{p})}\,d\mathbf{p}=\mathbb{E}_{\mathbf{p}\sim q(\mathbf{p})}\left[\frac{r(\mathbf{p})^{2}}{q(\mathbf{p})}\right].$ (12)
@@
$L\approx\hat{L}=\frac{1}{N}\sum_{i=1}^{N}r(i)^{2}\,\frac{1}{q(i)}.$ (13)
@@
$\frac{\partial\hat{L}}{\partial\boldsymbol{\beta}_{j}}=\frac{2}{N}\sum_{i=1}^{N}\frac{\partial r_{i}}{\partial\beta_{j}}r_{i}\,\frac{1}{q(i)}$ (14)
This gradient estimates the right-hand side of the normal equation given in Eq. 8 since $\frac{\partial r_{i}}{\partial\beta_{j}}=J_{ij}$.
@@
4.5 Dynamic Learning Rate Scheduler
When second-order optimizers are applied to large-scale systems, numerical errors can occur, causing the solution vector $\Delta\boldsymbol{\beta}$ to overshoot the true loss landscape. To determine the optimal learning rate, line search algorithms or trust region methods [2, 20, 37, 12, 42, 8] are usually employed. However, these methods require additional forward or backward passes, increasing the computational overhead and slowing down the overall algorithm. Rather than relying on these methods, we estimate the learning rate by constraining the maximum update of color values. In 3DGS, the color parameters range between $\approx-1.77$ and $\approx 1.77$ due to the level 0 spherical harmonic coefficient. This gives us a natural bound for the color parameters. We trust the update direction $\Delta\boldsymbol{\beta}$, as long as the resulting color change does not exceed 1. If the change surpasses this threshold, we scale $\Delta\boldsymbol{\beta}$ so that the maximum change remains 1. Notably, this scaling is applied uniformly across all parameters, regardless of their type (e.g., opacity, color, mean, scale, or rotation). The effectiveness of this learning rate heuristic is demonstrated in Tab. 4, and examples of assigned learning rates are given in the Supplementary.
5 Results
Datasets and Metrics. We conduct our main experiments on the synthetic NeRF [36] and real-world Mip-NeRF360 indoor datasets [4]. In the synthetic dataset, we position 10,00010{,}000 Gaussians at random locations with random colors inside of the cube encapsulating the object. In the real-world scenes, the Gaussians are initialized with structure-from-motion point clouds as in 3DGS [22]. For all experiments, we use the default training and test splits. We report the test set performance based on structural similarity index (SSIM) [52], learned perceptual image patch similarity score (LPIPS) [57], and peak signal-to-noise ratio (PSNR).
Baselines. In the synthetic dataset, we compare our optimizer against Adam [24], RMSprop [51], and SGD with momentum [46], using their respective implementations in PyTorch. The gradients are computed using either the vanilla 3DGS or Taming 3DGS [31], which enhances the efficiency of the backward computation by storing the gradient state at every $32^{\text{nd}}$ splat.
In the real-world scenes, we compare against vanilla 3DGS [22] and 3DGS-LM [20]. For comparisons against 3DGS-LM, we contacted the authors [^1] At the time of our submission, the code for 3DGS-LM had not yet been released., who kindly provided PSNR scores, runtime, and average memory usage across several datasets. In addition, they provided results obtained without the Adam optimizer, which are denoted in the tables as 3DGS-LM (w/o Adam). First-order baselines are run for $10,000$ iterations, and we performed a hyperparameter search to find the best learning rates.
Implementation Details. We conducted synthetic dataset experiments on a NVIDIA Tesla A40 GPU and real-world experiments on a NVIDIA Tesla A100 GPU. Our method and the baselines optimize the mean squared error (MSE) loss. We set the sampled pixels per tile ($N$) to 32 in each experiment. We use a fixed damping parameter $\lambda$ of $0.1$ for the synthetic NeRF scenes and Mip-NeRF360 dataset. For the synthetic dataset, the batch size is fixed at $8$; for real-world datasets, it is initialized at $16$ and increased to $32$ after the 50th iteration. The maximum number of conjugate gradient iterations starts at $3$ for synthetic datasets and at $5$ for real-world datasets, rising to $8$ after the 50th iteration.
[…]
3DGS 0.854 28.21 252 5.81
3DGS-LM (w/o Adam) - 26.72 1186 $\approx$ 50
3DGS-LM - 28.31 398 $\approx$ 50
LM-RS (Ours) 0.859 28.59 190 14.75
@@ 5.1 Comparison
Please refer to the Supplementary for the results using SSIM loss, as well as quantitative results in Tanks & Temples [25] and DeepBlending scenes [17].
@@ 5.2 Ablation Studies
Dynamic Learning Rate Scheduler. We compare our learning rate heuristic introduced in Sec. 4.5 against uniform learning rate, Armijo line search [2, 42], and a grid search method in Tab. 4. The grid search method selects the learning rate at each iteration by identifying the value that results in the greatest reduction in the objective function. Although this method is stable, the learning rates obtained are too pessimistic, and it cannot reach the quality of our heuristic. Armijo line search improves the grid search method by picking the highest learning rate that results in sufficient reduction of the loss. While this approach has been successfully integrated as an improvement in deterministic second-order optimizers, our findings indicate that its effectiveness is limited in stochastic settings, such as ours, where optimization relies on multiple approximations. We also found that the uniform learning rate can diverge (LR=0.1) or behave suboptimally (LR=0.07) compared to our scheduler.
@@
6 Limitations
We use only a diagonally approximated SSIM loss for performance reasons, which may affect the convergence behavior. Future work could explore more efficient implementations enabling the use of full SSIM loss. In addition, our matrix-free implementation requires more intermediate vectors than Adam, leading $3\times$ higher memory usage.
7 Conclusion
By leveraging the inherent sparsity of the Jacobian matrix and integrating a GPU-parallelized conjugate gradient solver, our method significantly reduces the computational overhead. Our novel view and pixel-wise sampling further enhance efficiency, enabling rapid convergence by decreasing the per-step overhead. Additionally, our dynamic learning rate scheduler eliminates the need for costly line search procedures, further accelerating training. Our approach achieves up to $5\times$ speedup over Adam, particularly excelling in scenarios with low number of Gaussians. Overall, our results highlight the potential of second-order methods in accelerating optimization for 3D Gaussian Splatting. We anticipate that future work will refine these techniques, particularly by incorporating additional loss terms without incurring runtime overhead, further advancing the efficiency and quality of Gaussian-based scene representations.
@@ Appendix A Details of CUDA Implementation
Algorithm 3 Forward Differentiation - LM-RS
@@
We follow the parallelization pattern applied in 3DGS [22], dividing images into $16\times 16$ blocks, and launching a thread for each pixel. Our main difference from 3DGS is that our block size is significantly less than $16\times 16=256$, due to pixel sampling. In our implementation, we only support sample sizes that are multiples of $32$, which allows us to do fast warp-level reductions when needed. Now, we describe our CUDA kernels in detail:
Jv Kernel. This kernel is similar to the backward pass of 3DGS and is executed in two kernel calls using backward mode differentiation. The first kernel computes and forwards the derivatives associated with splatting operations to the Gaussians, while the second kernel computes the derivatives related to Gaussian preprocessing. The main difference with 3DGS is that the number of threads per block is significantly reduced due to sampling, enabling us to perform faster computations. We use warp-level reductions to collect the partial dot products when needed.
Jv Kernel. This kernel implements forward‐mode differentiation via the dual‐number data structure. We overload each operator so that, in a single pass, it performs both the real‐value computation and its derivative. This approach is more efficient than using backward-mode differentiation when the Jacobian-vector product is needed [5].
Conceptually, this kernel works similarly to the forward pass of 3DGS, but enhanced to carry the derivative information. Therefore, unlike Jv kernel, it first carries out Gaussian preprocessing, then implements splatting. Since we should additionally carry the derivative information, the shared memory usage is doubled compared to the forward pass of 3DGS. For a comparison between the implementations of the forward pass of 3DGS and our forward-mode differentiation, see Alg. 2 and Alg. 3. As it can be observed from the algorithm, there is no need for warp reductions in this kernel.
@@ Appendix B More Results in Synthetic Dataset
We compare our optimizer against Adam [24], RMSprop [51], and SGD with momentum [46], using their respective implementations in PyTorch. The momentum term of SGD and second-moment factor of RMSprop is set to $0.99$. For SGD, we used the learning rates $0.16$ for the location, $0.2$ for color and $0.1$ for other parameters. We also apply a learning rate schedule for the mean value parameters, as suggested in 3DGS. The gradients are computed using either the vanilla 3DGS or Taming 3DGS [31]. All baselines are run for 10,00010{,}000 iterations. We provide the per-scene scores obtained in the NeRF synthetic dataset [36] in Table 6. Please also see the supplementary video, comparing Adam and our optimizer in real-time. Additionally, we provide visual results of competing methods in Fig. 6.
Results with SSIM Loss: We also conduct experiments using SSIM loss, which is approximated diagonally for performance reasons, as mentioned in the limitations section of the main paper. The usual SSIM loss propagates gradients through the local neighbors, which is hard to implement with the parallelization scheme of our $\mathbf{Diag(J^{\top}J)}$ kernel. Our diagonal approximation is similar that used in 3DGS-LM, which follows the derivations in [58]. In Table 7, we show that our optimizer still outperforms the competing methods even with the diagonal approximation. However, to compensate for this discrepancy, we ran our method for $300$ iterations instead of $200$ iterations. The weights are set to $1.0$ and $0.2$, for MSE and SSIM terms respectively. All other hyperparameters remain identical to those specified in the main paper.
@@ Appendix C More Results in Real-world Scenes
Results with SSIM Loss: As mentioned earlier, both 3DGS-LM and our implementation uses a diagonal approximation. In Tab. 11, we compare both methods and show that our optimizer converges faster than 3DGS-LM (w/o Adam). However, in this setting, the diagonal approximation becomes overly restrictive, and 3DGS-LM outperforms our method by leveraging the full SSIM loss in the initial steps, which the Adam optimizer conveniently handles.